# Strict area law entanglement versus chirality

### v1

@@ Abstract
Chirality is a property of a gapped phase of matter in two spatial dimensions that can be manifested through non-zero thermal or electrical Hall conductance. In this paper, we prove two no-go theorems that forbid such chirality for a quantum state in a finite dimensional local Hilbert space with strict area law entanglement entropies. As a crucial ingredient in the proofs, we introduce a new quantum information-theoretic primitive called instantaneous modular flow, which has many other potential applications.
@@ 2.  Preliminaries
(1) Setup: We shall consider a quantum many-body state $\left|\Psi\right\rangle$ in a tensor product Hilbert space with finite local dimension. We explicitly unpack these words as follows: The state $\left|\Psi\right\rangle$ describes a quantum many-body system on a two-dimensional lattice with local degrees of freedom living on the lattice sites. The Hilbert space $\mathcal{H}$ for $\left|\Psi\right\rangle$ is a tensor product of local Hilbert spaces, $\mathcal{H}=\otimes_{v}\mathcal{H}_{v}$, where $\mathcal{H}_{v}$ is the finite dimensional Hilbert space for the local degrees of freedom on the lattice site $v$ [^1] If the local degrees of freedom are fermionic, then the tensor product is $\mathbb{Z}_{2}$ graded.. Throughout this paper, we shall exclusively use $\left|\Psi\right\rangle$ to denote a quantum state under such a setup.
@@
In fact, the no-go theorems we shall discuss in the paper concern quantum states that satisfy a condition called bulk A1 introduced in [13], which is a relaxation of the strict area law condition. Explicitly, we say $\left|\Psi\right\rangle$ satisfies the bulk A1 if for any disk-like region with partition $BCD$ topologically equivalent to Fig. 1 (Left), the entanglement entropies from $\left|\Psi\right\rangle$ on these regions satisfy
@@
where $(S_{X})_{\left|\Psi\right\rangle}\equiv S(\rho_{X})$. With Eq. (2.1) one can check that a strict area law state satisfies bulk A1. On the other hand, A1 states include more than strict area law states, since A1 could still hold if there are additional contributions to Eq. (2.1) [^2] In general, bulk A1 allows any additional short-range entanglement contributions in the form of sum of local terms along $\partial A$ as discussed in [23]. Moreover, sharp corners of the region can make additional contributions to Eq. (2.1), which are consistent with A1. In fact, we show in [24] that such a corner contribution is a necessary consequence of $c_{-}\neq 0$. However, as a consequence of the no-go theorem 3.1, this violation of the exact area law cannot be the only one that results from $c_{-}\neq 0$ – since a corner contribution is consistent with A1, further violation is required.. An important fact about a state that satisfies bulk A1 is
@@
(3) Computation of chiral central charge $c_{-}$ and electric Hall conductance $\sigma_{xy}$: Previously, we have introduced chiral central charge as a property about the edge energy current. In fact, it is also a property of the bulk state. If $\left|\Psi\right\rangle$ under the setup discussed previously is a suitable representative state of a chiral gapped phase with non-vanishing thermal Hall conductance, one can compute $c_{-}$ using bulk reduced density matrices of $\left|\Psi\right\rangle$ through [^3] Similarly to the situation for topological entanglement entropy [25, 26], there are pathological states that violate the relation between $c_{-}$ and $J$ [27]. These examples also violate A1. Our result concerns only that $J\neq 0$ and does not directly rely on the relation between $c_{-}$ in the thermal Hall conductance and $J$. the modular commutator [16, 15]
@@
where $K_{\bullet}=-\ln\rho_{\bullet}$ is called modular Hamiltonian whose support indicated by the subscript is on region shown in Fig. 1 (Right). We list several properties of modular commutators in App. C.1. To compute electric Hall conductance, $\left|\Psi\right\rangle$ shall have an on-site $\textsf{U}(1)$ symmetry $U(t)=\prod_{v}U_{v}(t)$, i.e. $U(t)\left|\Psi\right\rangle\propto\left|\Psi\right\rangle$. $U_{v}(t)=e^{{\bf i}Q_{v}t}$ is generated by $Q_{v}$, which is supported on a lattice site $v$ and measures the amount of $\textsf{U}(1)$ charges on $v$. Then, the local charge operator on a region $X$ is defined as $Q_{X}=\sum_{v\in X}Q_{v}$. On such a $\textsf{U}(1)$ symmetric state $\left|\Psi\right\rangle$, it was proposed in [17] that the Hall conductance can be computed via
@@
using the same tripartite region shown in Fig. 1 (Right) [^4] We remark that to show $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ is a topological quantity, Ref. [17] assumed a so-called cluster property of the state. We discover that just the bulk A1 is enough. The proof of this fact is given in App. C.2..
@@ 3.  Statement of the no-go theorems
There is no quantum many-body state $\left|\Psi\right\rangle$ with on-site $\textsf{U}(1)$ symmetry in a tensor product Hilbert space with finite local dimension satisfying bulk A1 with non-zero $\sigma_{xy}$ from $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ using Eq. (2.5).
@@ 4.  Instantaneous modular flow
Definition. Consider a multi-partite quantum state $\left|\psi\right\rangle$ on a tensor product Hilbert space $\mathcal{H}=\otimes_{i}\mathcal{H}_{A_{i}}$, where $\mathcal{H}_{A_{i}}$ is the Hilbert space of a party $A_{i}$ of finite dimension. We will call a single party or a union of parties a region, although it may not be associated with a spatial region. We use $\overline{A}$ to denote the complement of the region $A$.
@@
We emphasize that the actual unitary operator that implements an instantaneous modular flow is made from the reduced density matrix from the current state, and is not a linear operator. This is the main difference between instantaneous modular flows and (ordinary) modular flows defined in the existing literature (e.g. [28, 29, 30, 31]). For example, to compute $\mathcal{I}_{B}(t)\mathcal{I}_{A}(s)\left|\psi\right\rangle$, first do a modular flow $\left|\psi^{\prime}\right\rangle=\mathcal{I}_{A}(s)\left|\psi\right\rangle=\rho_{A}^{{\bf i}s}\left|\psi\right\rangle$ using $\rho_{A}$ from $\left|\psi\right\rangle$. Then for $\mathcal{I}_{B}(t)$, use $\rho^{\prime}_{B}$ from the current state $\left|\psi^{\prime}\right\rangle$ to obtain $(\rho^{\prime}_{B})^{{\bf i}t}\left|\psi^{\prime}\right\rangle$ as the final result, which is in general different from $\rho_{B}^{{\bf i}t}\rho_{A}^{{\bf i}s}\left|\psi\right\rangle$ with both $\rho_{A},\rho_{B}$ from $\left|\psi\right\rangle$.
@@
Applications. One can apply instantaneous modular flows to various contexts. Here we discuss its application in gapped quantum phases of matter and quantum information. More applications are mentioned in Section 6.
@@
where $\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\equiv\left(\mathcal{I}_{AB}(x)\left|\Psi\right\rangle,\mathcal{I}_{BC}(y)\left|\Psi\right\rangle\right)$. Here we use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote state overlap: $(\left|\varphi\right\rangle,\left|\psi\right\rangle)\equiv\langle\varphi|\psi\rangle$. Similarly, one can compute $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ via
$\Sigma(A,B,C)_{\left|\Psi\right\rangle}=2{\rm Im\hskip 1.00006pt}\Big{[}\partial_{x}\partial^{2}_{y}\mathcal{F}^{A,B,C}_{\Sigma}(x,y)_{\left|\Psi\right\rangle}\Big{]}_{x,y=0}$ (4.5)
@@ 5.  Proofs of the no-go theorems
If a quantum state $\left|\Psi\right\rangle$ satisfies bulk A1, for a sufficiently large tripartite region $ABC$ shown in Fig. 1 (Right)
@@
The insight behind this statement is the following two points: [P1] With bulk A1, one can do Markov decomposition moves for the instantaneous modular flows on $\left|\Psi\right\rangle$. [P2] Instantaneous modular flows on $\left|\Psi\right\rangle$ preserve Markov conditions. Note that in showing Lemma 5.1 we only need a special case of P2, which we will prove explicitly. The proof of the general case is deferred to [32].
@@
In showing Eq. (5.9), we want to relate $\mathcal{I}_{AB}(x)\left|\Psi(t)\right\rangle$ and $\mathcal{I}_{BC}(y)\left|\Psi(t)\right\rangle$ to $\mathcal{I}_{AB}(x)\left|\Psi^{\prime}(t)\right\rangle$ and $\mathcal{I}_{AB}(y)\left|\Psi^{\prime}(t)\right\rangle$ by a unitary. Firstly, we can use the flipping move $\left|\Psi(t)\right\rangle=\mathcal{I}_{AB}(t)\left|\Psi\right\rangle=\mathcal{I}_{\overline{AB}}(t)\left|\Psi\right\rangle$. Secondly, with $\overline{AB}$ divided into small pieces as shown in Fig. 2, one can use Markov decomposition moves to write $\mathcal{I}_{\overline{AB}}(t)$ into a sequence of instantaneous modular flows supported on these small pieces. This enables us to use the commutation move to pull $\mathcal{I}_{AB}(x)$ and $\mathcal{I}_{BC}(y)$ through some of the instantaneous modular flows in the sequence, obtaining
@@
Proof sketch of Theorem 3.2: Roughly speaking, if there were a state $\left|\Psi\right\rangle$ with the properties described in Theorem 3.2, then one could “pump” the $\textsf{U}(1)$ charges indefinitely to a certain local region via a modular flow, such that the charge fluctuation in the region would be infinite. This is in contradiction to the setup that $Q_{v}$ has a bounded spectrum for all $v$. If the spectrum of $Q_{v}$ were unbounded, i.e. it had an infinite eigenvalue, then the unitary $U_{v}(t)$ would be ill-defined and so would be the $\textsf{U}(1)$ symmetry operator $U(t)=\prod_{v}U_{v}(t)$. The details are in App. E.
Remarks: Firstly, both no-go theorems are locally applicable. Even just on a local region777with sufficiently large, but still finite size $\left|\Psi\right\rangle$ can not satisfy strict area law with $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$. This is because the proofs only involve local entanglement properties of the quantum state: the strict area law is expressed with a local condition bulk A1, and $c_{-},\sigma_{xy}$ can be computed also locally from the state. Secondly, in the setup of $\left|\Psi\right\rangle$, it is the finite local Hilbert dimension that contradicts to bulk A1 and $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, while the tensor product structure can be relaxed. For example, as discussed in App. B, the no-go theorems are applicable for fermionic systems, where the tensor product is $\mathbb{Z}_{2}$ graded.
@@ 6.  Discussion
Our no-go theorems reveal that chirality, in the form of $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, constrains the local entanglement properties of a representative state of a chiral gapped phase. The no-go theorems and their proofs might give a hint on studying the subleading correction terms to the area law entanglement formula Eq. (2.1). For a representative of a chiral gapped phase in a finite-dimensional local Hilbert space, we expect that the bulk A1 is approximately satisfied with violations decreasing as the subsystem sizes are enlarged, consistent with the decay of the subleading corrections to Eq. (2.1). On such a state, it is expected that $J$ and $\Sigma$ compute $c_{-}$ and $\sigma_{xy}$ also with decreasing error for larger subsystem sizes. One example of such a case is studied in the numerical section of [22]. In fact, the techniques used in the proofs might be used to relate the violation of the Markov flow equation Eq. (4.6), which is a signature of the violation of bulk A1, to the deviations of $J$ and $\Sigma$ from the expected values. The detailed relation is left for future investigation.

### v2

@@ Abstract
Chirality is a property of a gapped phase of matter in two spatial dimensions that can be manifested through non-zero thermal or electrical Hall conductance. In this paper, we prove two no-go theorems that forbid such chirality for a quantum state in a finite dimensional local Hilbert space with strict area law entanglement entropies. We also show that the finite dimensional local Hilbert space condition can be relaxed to the condition that state has finite local entanglement entropies. As a crucial ingredient in the proofs, we introduce a new quantum information-theoretic primitive called instantaneous modular flow, which has many other potential applications.
@@ 2.  Preliminaries
(1) Setup: We shall consider a two-dimensional quantum many-body state $\left|\Psi\right\rangle$ in a tensor product Hilbert space with finite local dimension. We explicitly unpack these words as follows: The state $\left|\Psi\right\rangle$ describes a quantum many-body system on a two-dimensional lattice. The local degrees of freedom live on the lattice sites $v$ with finite dimensional Hilbert spaces $\mathcal{H}_{v}$. The global Hilbert space for $\left|\Psi\right\rangle$ is a tensor product $\mathcal{H}=\otimes_{v}\mathcal{H}_{v}$ [^1] If the local degrees of freedom are fermionic, then the tensor product is $\mathbb{Z}_{2}$ graded.. Throughout this paper, we shall exclusively use $\left|\Psi\right\rangle$ to denote a quantum state under such a setup.
@@
In fact, the no-go theorems we shall discuss later concern quantum states that satisfy a condition called bulk A1 introduced in [13], which relaxes the strict area law condition. Explicitly, we say $\left|\Psi\right\rangle$ satisfies the bulk A1 if for any disk-like region with partition $BCD$ topologically equivalent to Fig. 1 (Left), the entanglement entropies from $\left|\Psi\right\rangle$ on these regions satisfy
@@
where $(S_{X})_{\left|\Psi\right\rangle}\equiv S(\rho_{X})$. With Eq. (2.1) one can check that a strict area law state satisfies bulk A1. On the other hand, A1 states include more than strict area law states, since A1 could still hold if there are additional contributions to Eq. (2.1)  [^2] In general, bulk A1 allows any additional short-range entanglement contributions in the form of sum of local terms along $\partial A$ as discussed in [23]. An explicit state satisfying A1 but not strict area law can be obtained by starting with a state with strict area law and tensoring in EPR pairs $\sqrt{p_{uv}}\left|0_{u}0_{v}\right\rangle+\sqrt{1-p_{uv}}\left|1_{u}1_{v}\right\rangle$ of varying strength $p_{uv}$ along each link $<uv>$ of the lattice. The whole state satisfies A1 still, but the area law coefficient will vary in space according to the choices of $p$. Moreover, sharp corners of the region can make additional contributions to Eq. (2.1), which are consistent with A1. In fact, we show in [24] that such a corner contribution is a necessary consequence of $c_{-}\neq 0$. However, as a consequence of the no-go theorem 3.1, this violation of the exact area law cannot be the only one that results from $c_{-}\neq 0$ – since a corner contribution is consistent with A1, further violation is required.. An important fact about a state that satisfies bulk A1 is
@@
(3) Computation of chiral central charge $c_{-}$ and electric Hall conductance $\sigma_{xy}$: Previously, we have introduced chiral central charge as a property about the edge energy current. In fact, it is also a property of the bulk state. If $\left|\Psi\right\rangle$ is a suitable representative state of a chiral gapped phase with non-vanishing thermal Hall conductance, one can compute $c_{-}$ using bulk reduced density matrices of $\left|\Psi\right\rangle$ through [^3] Similarly to the situation for topological entanglement entropy [25, 26], there are pathological states where $J$ is not associated with a meaning of chiral central charge [27]. These examples also violate A1. Nevertheless, $J\neq 0$ is an indication of chirality of the state regardless whether it is related to the thermal Hall conductance, and our result indicates such a chirality is not compatible with strict area law state.
@@
where ${\bf i}=\sqrt{-1}$ and $K_{\bullet}=-\ln\rho_{\bullet}$ is called modular Hamiltonian whose support indicated by the subscript is on region shown in Fig. 1 (Right). To compute electric Hall conductance, $\left|\Psi\right\rangle$ shall have an on-site $\textsf{U}(1)$ symmetry $U(t)=\prod_{v}U_{v}(t)$, i.e. $U(t)\left|\Psi\right\rangle\propto\left|\Psi\right\rangle$. $U_{v}(t)=e^{{\bf i}Q_{v}t}$ is generated by $Q_{v}$, which is supported on a lattice site $v$ and measures the amount of $\textsf{U}(1)$ charges on $v$. Then, the local charge operator on a region $X$ is defined as $Q_{X}=\sum_{v\in X}Q_{v}$. On such a $\textsf{U}(1)$ symmetric state $\left|\Psi\right\rangle$, it was proposed in [17] that the Hall conductance can be computed via
@@
using the same tripartite region shown in Fig. 1 (Right). We list several properties of $J$ and $\Sigma$ in App. C [^4] We remark that to show $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ is a topological quantity, Ref. [17] assumed a so-called cluster property of the state. We discover that just the bulk A1 is enough. The proof of this fact is given in App. C.2..
@@ 3.  Statement of the no-go theorems
There is no quantum many-body state $\left|\Psi\right\rangle$ in a tensor product Hilbert space with finite local dimension satisfying bulk A1 and having an on-site $\textsf{U}(1)$ symmetry and non-zero $\sigma_{xy}$ from $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ using Eq. (2.5).
We remark that the condition of finite local Hilbert space dimension in both Theorems 3.1 and 3.2 can be relaxed into finite local entanglement entropies. We leave the discussion of the generalized version in App. F.
@@ 4.  Instantaneous modular flow
Definition. Consider a multi-partite quantum state $\left|\psi\right\rangle$ on a tensor product Hilbert space $\mathcal{H}=\otimes_{i}\mathcal{H}_{A_{i}}$, where $\mathcal{H}_{A_{i}}$ is the Hilbert space of a party $A_{i}$ of finite dimension.  [^5] We consider this setup for simplicity and it is enough for the purpose of this paper. It is possible to generalize the definition for a system in a Hilbert space without tensor product structure or finite local dimension, such as continuum quantum field theory. We will call a single party or a union of parties a region, although it may not be associated with a spatial region. We use $\overline{A}$ to denote the complement of the region $A$.
@@
We emphasize that the actual unitary operator that implements an instantaneous modular flow is made from the reduced density matrix from the current state. An instantaneous modular flow should be regarded as a nonlinear map that generates a certain state evolution. This is the main difference between instantaneous modular flows and (ordinary) modular flows defined in the existing literature (e.g. [28, 29, 30, 31]). For example, to compute $\mathcal{I}_{B}(t)\mathcal{I}_{A}(s)\left|\psi\right\rangle$, first do a modular flow $\left|\psi^{\prime}\right\rangle=\mathcal{I}_{A}(s)\left|\psi\right\rangle=\rho_{A}^{{\bf i}s}\left|\psi\right\rangle$ using $\rho_{A}$ from $\left|\psi\right\rangle$. Then for $\mathcal{I}_{B}(t)$, use $\rho^{\prime}_{B}$ from the current state $\left|\psi^{\prime}\right\rangle$ to obtain $(\rho^{\prime}_{B})^{{\bf i}t}\left|\psi^{\prime}\right\rangle$ as the final result, which is in general different from $\rho_{B}^{{\bf i}t}\rho_{A}^{{\bf i}s}\left|\psi\right\rangle$ with both $\rho_{A},\rho_{B}$ from $\left|\psi\right\rangle$. We remark that the focal point of the instantaneous modular flow is to consider their sequences and the resulting state evolutions, while the ordinary modular flow is usually regarded as a unitary evolution for operators in some algebra.
@@
Applications. One can apply instantaneous modular flows to various contexts. Here we discuss its application in gapped quantum phases of matter and quantum information. More applications are mentioned in Sec. 6.
@@
where $\mathcal{F}^{A,B,C}_{J}(x,y)_{\left|\Psi\right\rangle}\equiv\left(\mathcal{I}_{AB}(x)\left|\Psi\right\rangle,\mathcal{I}_{BC}(y)\left|\Psi\right\rangle\right)$ [^7] Here we use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote state overlap: $(\left|\varphi\right\rangle,\left|\psi\right\rangle)\equiv\langle\varphi|\psi\rangle$. Similarly, one can compute $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ via
$\Sigma(A,B,C)_{\left|\Psi\right\rangle}=-\mathrm{Re}\Big{[}\partial_{x}\partial^{2}_{y}\mathcal{F}^{A,B,C}_{\Sigma}(x,y)_{\left|\Psi\right\rangle}\Big{]}_{x,y=0}$ (4.5)
@@ 5.  Proofs of the no-go theorems
The proof of Theorem 3.2 follows a similar strategy. Roughly speaking, if there were a state $\left|\Psi\right\rangle$ with the properties described in Theorem 3.2, then one could increase the charge fluctuation of a local region infinitely via modular flow, such that the charge fluctuation in the region would be infinite. This is in contradiction to the fact that $Q_{v}$ has a bounded spectrum for all $v$.
Below we present the detailed proof of Theorem 3.1, featured with usages of instantaneous modular flows. We leave the analogous proof of Theorem 3.2 in App. E.
@@
If $\left|\Psi\right\rangle$ satisfies bulk A1, for a sufficiently large tripartite region $ABC$ shown in Fig. 1 (Right)
@@
Proof sketch of Lemma 5.1. The insight behind this statement is the following two points: [P1] With bulk A1, one can do Markov decomposition moves for the instantaneous modular flows on $\left|\Psi\right\rangle$. [P2] Instantaneous modular flows on $\left|\Psi\right\rangle$ preserve Markov conditions. Note that in showing Lemma 5.1 we only need a special case of P2, which we will prove explicitly. The proof of the general case is deferred to [32].
@@
In showing Eq. (5.9), we want to relate $\mathcal{I}_{AB}(x)\left|\Psi(t)\right\rangle$ and $\mathcal{I}_{BC}(y)\left|\Psi(t)\right\rangle$ to $\mathcal{I}_{AB}(x)\left|\Psi^{\prime}(t)\right\rangle$ and $\mathcal{I}_{BC}(y)\left|\Psi^{\prime}(t)\right\rangle$ by a unitary. Firstly, we can use the flipping move $\left|\Psi(t)\right\rangle=\mathcal{I}_{AB}(t)\left|\Psi\right\rangle=\mathcal{I}_{\overline{AB}}(t)\left|\Psi\right\rangle$. Secondly, with $\overline{AB}$ divided into small pieces as shown in Fig. 2, one can use Markov decomposition moves to write $\mathcal{I}_{\overline{AB}}(t)$ into a sequence of instantaneous modular flows supported on these small pieces. This enables us to use the commutation move to pull $\mathcal{I}_{AB}(x)$ and $\mathcal{I}_{BC}(y)$ through some of the instantaneous modular flows in the sequence, obtaining
@@
Remarks: Firstly, both no-go theorems are locally applicable. Even just on a local region999with sufficiently large, but still finite size $\left|\Psi\right\rangle$ can not satisfy strict area law with $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$. This is because the proofs only involve local entanglement properties of the quantum state: the strict area law is expressed with a local condition bulk A1, and $c_{-},\sigma_{xy}$ can be computed also locally from the state. Secondly, in the setup of $\left|\Psi\right\rangle$, it is the finite local Hilbert dimension that contradicts to bulk A1 and $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, while the tensor product structure can be relaxed. For example, as discussed in App. B, the no-go theorems are applicable for fermionic systems, where the tensor product is $\mathbb{Z}_{2}$ graded.
@@ 6.  Discussion
Our no-go theorems reveal that chirality, in the form of $c_{-}\neq 0$ or $\sigma_{xy}\neq 0$, constrains the local entanglement properties of a representative state of a chiral gapped phase [^10] We comment that one might hope to combine the recent result of [33], which says that any state satisfying A0 and A1 has a commuting projector parent Hamiltonian, with the Hamiltonian no-go theorems of [6, 7, 8], to provide a proof of a related result.. Loosely speaking, our no-go theorems indicate that it is impossible to construct a state in a Hilbert space with finite local dimensions (or alternatively, with finite local entanglement entropies) that represent a zero correlation length fixed point of a chiral gapped phase. [^11] This also aligns with various difficulties when people attempt to construct a chiral gapped state, such as using PEPS [Hasik:2022, Wahl:2013rha] or MERA [Chu:2023ven]. We leave the detailed connections to be investigated in the future. Moreover, the no-go theorems and their proofs might give a hint on studying the subleading correction terms to the area law entanglement formula Eq. (2.1). For a representative of a chiral gapped phase in a finite-dimensional local Hilbert space, we expect that the bulk A1 is approximately satisfied with violations decreasing as the subsystem sizes are enlarged, consistent with the decay of the subleading corrections to Eq. (2.1). On such a state, it is expected that $J$ and $\Sigma$ compute $c_{-}$ and $\sigma_{xy}$ also with decreasing error for larger subsystem sizes. One example of such a case is studied in the numerical section of [22]. In fact, the techniques used in the proofs might be used to relate the violation of the Markov flow equation Eq. (4.6), which is a signature of the violation of bulk A1, to the deviations of $J$ and $\Sigma$ from the expected values. The detailed relation is left for future investigation.